# Bayesian Controlled FDR Variable Selection via Knockoffs

### v2

@@ 3 Simulation study
We evaluate the performance of the proposed model through a comprehensive simulation study and compare results with those of the classical model-X knockoff filter and state-of-the-art variable selection methods, including lasso regression (Tibshirani,, 1996) and models that use discrete spike-and-slab priors (Mitchell and Beauchamp,, 1988).
@@ 3.1 Data generation
We generated data from a few distinct scenarios. In Scenario $1$, we generated $p=30$ covariates independently sampled from a standard Gaussian distribution. Out of the $30$ covariates, we assumed that $6$ of them have a linear effect on the response variable $y$. The regression coefficients associated with these active covariates were randomly selected from the set {±$\{\pm 0.5,\pm 1,\pm 1.5\}$ to introduce variation in the strength of the effects. We then simulated the response variable from the linear equation $y_{i}=x_{i}\beta+\epsilon_{i}$, with $\epsilon_{i}\sim N(0,\sigma_{\epsilon})$ and $\sigma_{\epsilon}=1$, for $i=1,\ldots,n$. This relatively simple scenario enables us to assess the capability of our model to recover the set of active variables when the graphical structure learning is, effectively, a source of noise in the model. In Scenario $2$, we introduce a more realistic setting by incorporating a correlation structure among the predictors. More specifically, the $30$ covariates were generated from a Gaussian graphical model, where the underlying graph structure is characterized by a sparse precision matrix. The underlying graph structure is shown in Figure 1. Similarly to Scenario $1$, we selected $6$ out of the $30$ variables to be active in explaining the response $y$ and randomly sampled their coefficient values from the set {±$\{\pm 0.5,\pm 1\}$. In Scenario $3$, data are simulated following the setting of Li and Li, (2008) and Peterson et al., (2016). Specifically, we generated the covariates from a graph with 40 hubs, each with 5 connected nodes. This resulted in $240$ covariates with a sparse graphical model of $200$ edges in total. We set as active variables those in the first $4$ hubs, plus their connected edges, resulting in a total of $24$ non-zero coefficients. In particular the regression coefficients of the 4 hubs are $5$, $-5$, $3$ and $-3$, respectively, while the coefficients of the associated nodes are the same but divided by $\sqrt{10}$. The covariates were then generated from a Gaussian distribution with zero mean and covariance matrix $\bm{\Sigma}$ with diagonal elements equal to $1$ and covariances between connected variables equal to $0.7$. As in the previous scenarios, the response variable was simulated from a linear Gaussian model with the error variance equal to $\sum_{j}\beta_{j}^{2}/4$. In each scenario we set the sample size at $n=200$ and considered $100$ replicated data sets.
Figure 1: Simulation study: Scale-free graph structure under Gaussian graphical model in Scenario $2$. White nodes correspond to the noise variables, and blue nodes to the active ones. A stronger blue color corresponds to a larger effect.
@@ 3.2 Parameter settings
When fitting our model, we need to specify the hyperparameters of the shrinkage inducing prior (10) on the precision matrix of the covariates’ graph, the spike-and-slab priors (12) on the regression coefficients and the Ising prior (13) on the selection indicators. In all analyses, for prior (10) on the precision matrix, following the recommendations in Wang, (2015), we used $v_{0}=0.01^{2}$ and $v_{1}=hv_{0}$ with $h=100$. We set the prior probability of edge selection $\xi$ to $0.01$ and the exponential parameter $\theta$ for the elements on the diagonal to $2$. To specify vaguely informative priors, we set $h_{\beta}=1$ in the joint prior (12) on the regression coefficients, and $\alpha_{\sigma}=\beta_{\sigma}=2$ for the prior (14) on the error variance. Furthermore, we centered covariates and response variables, ensuring that the intercept term in the model is zero. Smith, (1973) showed that, in the Bayesian framework, setting the intercept term to zero is equivalent to placing a Gaussian prior with infinite variance on the intercept parameter. Finally, for the Ising prior (13) on $\gamma$, we set $a=b=0.5$. This specification corresponds to a prior inclusion probability of $0.622$ on the original variable or its knockoff, when the graph is empty. When experimenting with this prior, we found non-sparse specifications to work well, as the Bayesian knockoff framework allows for further control of the FDR by excluding inactive covariates. We provide sensitivity analyses to these prior choices in Section 3.5.
When initializing the MCMC algorithm, we set the selection indicators $\gamma$ to 00 and assumed independent covariates with $\bm{\Omega}=\bm{I}$. The results reported below in Section 3.3 are obtained by running MCMC chains with $8,000$ iterations for burn-in, followed by $8,000$ iterations for inference. For each chain, we assessed convergence by using the Geweke’s diagnostic test (Geweke,, 1992) to check for signs of non-convergence. For example, the average $z$-scores from the test for $\{(\beta_{j},\widetilde{\beta}_{j})\}_{j=1}^{p}$ were $0.03$, $0.01$ and $-0.05$ for Scenario $1$, Scenario $2$ and Scenario $3$, respectively. This suggests that the algorithm has run for enough iterations.
@@ 3.3 Results and comparisons
Let us first examine a single data set to gain insight into how the proposed model works. Specifically, we focus on a randomly selected data set generated under Scenario 3. The top plot of Figure 2 shows the estimated BFDR and the upper bound for $\mathbb{P}[r_{j}=0\mid\bm{D}]$, for $j=1,\ldots p$ and covariates sorted by the BFDR values. Our method correctly identifies 18 of the 24 true predictors with 6 false negatives, for a $q=0.1$ threshold on the BFDR. The bottom plot of Figure 2 reports the PPIs of all covariates, calculated by fitting a traditional Bayesian linear regression model with discrete spike-and-slab priors and a Beta hyperprior on the inclusion probability of the latent indicators (George and McCulloch,, 1997; Brown et al.,, 1998; Vannucci,, 2021). This method selects 21 of the 24 true predictors with 44 false positives, for a $0.5$ threshold on the PPIs. The large number of false positives is due to the fact that the presence of collinearity in the data greatly affects the PPIs values, resulting in inflated values for inactive covariates that are correlated to the active ones. This simple example highlights the inability of clearly dividing active and inactive variables based on the posterior probability values. The results are confirmed in the comparison study below, that shows performance values averaged over replicated datasets. This empirical comparison of our Bayesian Knockoff with a traditional spike and slab regression confirms the theoretical results in Theorem 1, where the upper bound is unaffected by the covariates’ dependency structure.
@@
Next, we report results of our method averaged over 100 replicated datasets and perform a comparison of the performances with other procedures for variable selection, with and without FDR control. Among these procedures, we include the classical model-X knockoff filter proposed by Barber and Candès, (2019). In particular, we consider two versions of the model-X knockoff filter, one with the true covariance matrix assumed known, and one with an estimate of $\bm{\Sigma}$ obtained via the graphical lasso estimator of Friedman et al., (2008). Indeed, Barber et al., (2020) proved that, under certain conditions, the graphical lasso can be used to preserve the FDR guarantees of knockoff filters. For both our model and the classical knockoff filter, we fix the FDR threshold at $q=0.1$. The knockoffs were generated using the R package knockoffs, and the covariance matrix was estimated using the glasso package. In addition to the model-X knockoff filter, our comparison includes other state-of-art methods for variable selection. In particular, we include the Lasso regression (Tibshirani,, 1996) and a Bayesian linear regression model with discrete spike-and-slab prior and a Beta hyperprior on the inclusion probability, which we implemented ourselves. We also considered the joint edge and regression estimation model proposed by Peterson et al., (2016), which we implemented ourselves as a version of our proposed model without knockoffs latent variables. The regularizing parameter for lasso regression was determined by minimizing the cross-validated error using the R package glmnet. We note that this method does not employ any FDR correction. Variable selection for the Bayesian and joint estimation methods was achieved by thresholding the marginal posterior probabilities of inclusion (PPIs) at $0.5$, following the median model approach (Barbieri and Berger,, 2004). We also considered the BFDR approach proposed by Newton et al., (2004).
To compare performance, in Table 1 we report averages and standard deviations across $100$ replicates of several evaluation metrics. In particular, we consider FDR, True Positve Rate (TPR), Matthew Correlation Coefficient (MCC), and F1-score. Let TP be the number of true positives, TN the number of true negatives, FP the number of false positives, and FN the number of false negatives. We can then define the performance metrics used as
@@
Results in Table 1 indicate that the proposed Bayesian method succeeds in keeping the FDR below the given threshold in every scenario, while achieving strong performances in TPR compared to the other methods, as well as MCC and F1 scores. It is worth noting that the classical knockoff filter exhibits low TPR in Scenarios $1$ and $2$, where the magnitude of the true coefficients is rather small. In other settings, not reported here, where the signal was strong and the sample size larger, we noticed improved performances for all methods, including classical knockoffs. Further, the values of the standard deviations for the TPR of the knockoff filter in the first two scenarios are higher than those of the other methods. This happens because, across the $100$ simulated data sets, the knockoff filter TPR is often zero and sometimes quite high. Finally, we note that classical knockoff has a slightly higher FDR than the threshold $0.1$ when the covariates are not independent and the covariance matrix is estimated using graphical lasso. As shown in Barber et al., (2020), the classical knockoff filter does not control FDR exactly below $q$ when the covariance matrix used to build the knockoffs is not the true one but an estimate. In our case, the graphical lasso tends to produce a sparser than needed graph. This behavior also supports our choice of setting the parameter of the priors in the proposed method to obtain a generally dense graph to be used to generate the knockoffs. Results for the third scenario confirm that the proposed Bayesian Knockoff filter is also able to outperform the state-of-art methods in settings with larger dimensionality. As a matter of fact, it successfully controls FDR under $0.1$, unlike the lasso regression. Moreover, the TPR is on average comparable to the one of the classical knockoff filter, but it is more stable as it has a much lower standard deviation.
Table 1: Simulation study: FDR, TPR, MCC and F1-score, for the scenarios described in Section 3.1. Compared methods are the proposed model (BayesKnock) classical Knockoff filter with exact (Knockoff (Exact)) and estimated (Knockoff (Glasso)) covariance matrix, Lasso, spike-and-slab regression (Spike-Slab) and the joint model proposed by Peterson et al., (2016) (Joint). Values reported are averages over $100$ replicated data sets, with standard deviations in parentheses.
@@
As for the comparison with the state-of-art methods for variable selection, we can see that Lasso has a high TPR in every scenario but also a high FDR. Indeed, Meinshausen and Bühlmann, (2006) noted that the lasso estimator tends to include too many inactive predictors when the penalization parameter is chosen to have low prediction error, as we did in our simulations. Compared with Bayesian variable selection via spike-and-slab prior, our model offers comparable results in the first two scenarios, where the covariates are independent, with almost perfect results. In the third scenario, with dependent covariates, spike-and-slab loses the control of FDR, unlike our method, even though it manages to find more relevant predictors than our method. In Table 1 we do not show results for the BFDR approach proposed by Newton et al., (2004). Indeed, BFDR selected exactly the same variables as spike-and-slab in the first two scenarios. In the third scenario, however, the PPIs were hovering around $0.5$ and the BFDR did not select any variable. As pointed out in the introduction section, Bayes FDR selection procedures rely on the estimates of the marginal PPIs, as obtained from the MCMC output, while our procedure employs an upper bound on the probabilities of non-inclusion. Therefore, our model appears to achieve good operating characteristics and to control the FDR even when the estimates of the marginal posterior probability of inclusion fail to select the true relevant variables. This is evident when comparing our method with the model proposed by Peterson et al., (2016) (Joint). In scenarios 1 and 2 the two methods have the same performances, while in scenario 3, even though MCC and F1 have similar values, the addition of the knockoffs allows to keep the FDR under the desired level, while the joint method has higher FDR and therefore not an explicit control over it.
@@
our model achieved very solid performances for graph selection across the scenarios. For example, the average Frobenius norm was $1.404$, $1.415$ and $9.425$ and the F1-score was $0.869$, $0.865$ and $0.99$ in Scenarios $1$, $2$ and $3$, respectively. Note that the Frobenius norm is much higher in Scenario $3$ due to the different magnitude in the correlation between variables. However, as we can see from the F1-score, the graph recovery still works.
3.4 Simulation of a survival outcome from an accelerated failure time model
We also show a simulation for a regression model with a survival outcome, specifically an accelerated failure time (AFT) model, with censored observations. To use the proposed Bayesian knockoff filter, we adapt the stochastic search MCMC procedure of Sha et al., (2006), that makes use of data augmentation techniques (Tanner and Wong,, 1987). Let $T_{i}$ represent the time-to-event for the $i$-th subject and $c_{i}$ the censoring time. AFT models assume a multiplicative effect on the survival times of the general form $\log(T_{i})=X_{i}\beta+\epsilon_{i}$, where the errors $\epsilon_{i}$ may have several distributions. Here, we assume the errors to be i.i.d. Gaussian variables. To allow for censored observations, we introduce a latent variable $y_{i}$ such that
$\begin{cases}y_{i}=\log(T^{*}_{i})&\text{if }\kappa_{i}=1\\ y_{i}>\log(T^{*}_{i})&\text{if }\kappa_{i}=0\\ \end{cases}.$
where $T^{*}_{i}$ is defined as $\min(T_{i},c_{i})$ and $\kappa_{i}=\mathbbm{1}_{T_{i}\leq c_{i}}$ is the censoring indicator. Here, we assume that the latent variable $y_{i}$ given the covariates $X$ follows the model described in Equation (9) with $p(Y\mid X,U)$ as in Equation (8), with the same prior specification. We generate the covariates following the scheme employed in Scenario $3$ above, and then generate the log of the survival times from the linear AFT model. We allow for $25$ of the subjects to be censored. Table 2 shows performances of our proposed method, compared to the spike-and-slab method for AFT models of Sha et al., (2004) and to the same AFT model without knockoff latent variable, as also implemented by Peterson et al., (2016). These results are consistent with those for Scenario 3 in the previous simulation and also show that the proposed procedure suffers the least from the presence of censored data.
Table 2: Simulation study: FDR, TPR, MCC and F1-score, for the AFT models. Compared methods are the proposed model (BayesKnock), spike-and-slab regression (Spike-Slab) and the joint model proposed by Peterson et al., (2016) (Joint). Values reported are averages over $100$ replicated data sets, with standard deviations in parentheses.

### v3

@@ 3 Simulation study
We evaluate the performance of the proposed model through a comprehensive simulation study and compare results with those of the classical model-X knockoff filter and state-of-the-art variable selection methods, including lasso regression (Tibshirani, 1996) and models that use discrete spike-and-slab priors (Mitchell and Beauchamp, 1988).
@@ 3.1 Data generation
We generated data from a few distinct scenarios. In Scenario $1$, we generated $p=30$ covariates independently sampled from a standard Gaussian distribution. Out of the $30$ covariates, we assumed that $6$ of them have a linear effect on the response variable $y$. The regression coefficients associated with these active covariates were randomly selected from the set $\{\pm 0.5,\pm 1,\pm 1.5\}$ to introduce variation in the strength of the effects. We then simulated the response variable from the linear equation $y_{i}=x_{i}\beta+\epsilon_{i}$, with $\epsilon_{i}\sim N(0,\sigma_{\epsilon})$ and $\sigma_{\epsilon}=1$, for $i=1,\ldots,n$. This relatively simple scenario enables us to assess the capability of our model to recover the set of active variables when the graphical structure learning is, effectively, a source of noise in the model. In Scenario $2$, we introduce a more realistic setting by incorporating a correlation structure among the predictors. More specifically, the $30$ covariates were generated from a Gaussian graphical model, where the underlying graph structure is characterized by a sparse precision matrix. The underlying graph structure is shown in Figure 1. Similarly to Scenario $1$, we selected $6$ out of the $30$ variables to be active in explaining the response $y$ and randomly sampled their coefficient values from the set $\{\pm 0.5,\pm 1\}$. In Scenario $3$, data are simulated following the setting of Li and Li (2008) and Peterson et al. (2016). Specifically, we generated the covariates from a graph with 40 hubs, each with 5 connected nodes. This resulted in $240$ covariates with a sparse graphical model of $200$ edges in total. We set as active variables those in the first $4$ hubs, plus their connected edges, resulting in a total of $24$ non-zero coefficients. In particular the regression coefficients of the 4 hubs are $5$, 5-5, 33 and 3-3, respectively, while the coefficients of the associated nodes are the same but divided by $\sqrt{10}$. The covariates were then generated from a Gaussian distribution with zero mean and covariance matrix $\bm{\Sigma}$ with diagonal elements equal to $1$ and covariances between connected variables equal to $0.7$. As in the previous scenarios, the response variable was simulated from a linear Gaussian model with the error variance equal to $\sum_{j}\beta_{j}^{2}/4$. In each scenario we set the sample size at $n=200$ and considered $100$ replicated data sets.
Figure 1: Simulation study: Scale-free graph structure under Gaussian graphical model in Scenario 22. White nodes correspond to the noise variables, and black nodes to the active ones. A stronger black color corresponds to a larger effect.
@@ 3.2 Parameter settings
When fitting our model, we need to specify the hyperparameters of the shrinkage inducing prior (10) on the precision matrix of the covariates’ graph, the spike-and-slab priors (12) on the regression coefficients and the Ising prior (14) on the selection indicators. In all analyses, for prior (10) on the precision matrix, following the recommendations in Wang (2015), we used $v_{0}=0.01^{2}$ and $v_{1}=hv_{0}$ with $h=100$. We set the prior probability of edge selection $\xi$ to $0.01$ and the exponential parameter $\theta$ for the elements on the diagonal to $2$. To specify vaguely informative priors, we set $h_{\beta}=1$ in the joint prior (12) on the regression coefficients, and $\alpha_{\sigma}=\beta_{\sigma}=2$ for the prior (15) on the error variance. Furthermore, we centered covariates and response variables, ensuring that the intercept term in the model is zero. Smith (1973) showed that, in the Bayesian framework, setting the intercept term to zero is equivalent to placing a Gaussian prior with infinite variance on the intercept parameter. Finally, for the Ising prior (14) on $\gamma$, we set $a=b=0.5$. This specification corresponds to a prior inclusion probability of $0.622$ on the original variable or its knockoff, when the graph is empty. When experimenting with this prior, we found non-sparse specifications to work well, as the Bayesian knockoff framework allows for further control of the FDR by excluding inactive covariates. We provide sensitivity analyses to these prior choices in Section 3.4.
When initializing the MCMC algorithm, we set the selection indicators $\gamma$ to 0 and assumed independent covariates with $\bm{\Omega}=\bm{I}$. The results reported below in Section 3.3 are obtained by running MCMC chains with $8,000$ iterations for burn-in, followed by $8,000$ iterations for inference. For each chain, we assessed convergence by using the Geweke’s diagnostic test (Geweke, 1992) to check for signs of non-convergence. For example, the average $z$-scores from the test for $\{(\beta_{j},\widetilde{\beta}_{j})\}_{j=1}^{p}$ were $0.03$, $0.01$ and 0.05-0.05 for Scenario $1$, Scenario $2$ and Scenario $3$, respectively. This suggests that the algorithm has run for enough iterations.
@@ 3.3 Results and comparisons
Let us first examine a single data set to gain insight into how the proposed model works. Specifically, we focus on a randomly selected data set generated under Scenario 3. The top plot of Figure 2 shows the estimated BFDR and the upper bound for $\mathbb{P}[r_{j}=0\mid\bm{D}]$, for $j=1,\ldots p$ and covariates sorted by the BFDR values. Our method correctly identifies 18 of the 24 true predictors with 6 false negatives, for a $q=0.1$ threshold on the BFDR. The bottom plot of Figure 2 reports the PPIs of all covariates, calculated by fitting a traditional Bayesian linear regression model with discrete spike-and-slab priors and a Beta hyperprior on the inclusion probability of the latent indicators (George and McCulloch, 1997; Brown et al., 1998; Vannucci, 2021). This method selects 21 of the 24 true predictors with 44 false positives, for a $0.5$ threshold on the PPIs. The large number of false positives is due to the fact that the presence of collinearity in the data greatly affects the PPIs values, resulting in inflated values for inactive covariates that are correlated to the active ones. This simple example highlights the inability of clearly dividing active and inactive variables based on the posterior probability values. The results are confirmed in the comparison study below, that shows performance values averaged over replicated datasets. This empirical comparison of our Bayesian Knockoff with a traditional spike and slab regression confirms the theoretical results in Proposition 1, where the upper bound is unaffected by the covariates’ dependency structure.
@@
Next, we report results of our method averaged over 100 replicated datasets and perform a comparison of the performances with other procedures for variable selection, with and without FDR control. Among these procedures, we include the classical model-X knockoff filter proposed by Barber and Candès (2019). In particular, we consider two versions of the model-X knockoff filter, one with the true covariance matrix assumed known, and one with an estimate of $\bm{\Sigma}$ obtained via the graphical lasso estimator of Friedman et al. (2008). Indeed, Barber et al. (2020) proved that, under certain conditions, the graphical lasso can be used to preserve the FDR guarantees of knockoff filters. For both our model and the classical knockoff filter, we fix the FDR threshold at $q=0.1$. The knockoffs were generated using the R package knockoffs, and the covariance matrix was estimated using the glasso package. In addition to the model-X knockoff filter, our comparison includes other state-of-art methods for variable selection. In particular, we include the Lasso regression (Tibshirani, 1996) and a Bayesian linear regression model with discrete spike-and-slab prior and a Beta hyperprior on the inclusion probability, which we implemented ourselves. We also considered the joint edge and regression estimation model proposed by Peterson et al. (2016), which we implemented ourselves as a version of our proposed model latent knockoffs layer and a Bayesian variable selection model with non-local spike and slab proposed by Johnson and Rossell (2012) and implemented in the R package mombf. The regularizing parameter for lasso regression was determined as the largest value such that the cross-validation error is within one standard error of the minimum, using the R package glmnet. We note that this method does not employ any FDR correction. Variable selection for the Bayesian methods was achieved by thresholding the marginal posterior probabilities of inclusion (PPIs) at $0.5$, following the median model approach (Barbieri and Berger, 2004). We also considered the BFDR approach proposed by Newton et al. (2004).
To compare performance, in Table 1 we report averages and standard deviations across $100$ replicates of several evaluation metrics. In particular, we consider FDR, True Positive Rate (TPR), Matthews Correlation Coefficient (MCC), and F1-score. Let TP be the number of true positives, TN the number of true negatives, FP the number of false positives, and FN the number of false negatives. We can then define the performance metrics used as
@@
Results in Table 1 indicate that the proposed Bayesian method succeeds in keeping the FDR below the given threshold in every scenario, while achieving strong performances in TPR compared to the other methods, as well as MCC and F1 scores. It is worth noting that the classical knockoff filter exhibits low TPR in Scenarios $1$ and $2$, where the magnitude of the true coefficients is rather small. In other settings, not reported here, where the signal was strong and the sample size larger, we noticed improved performances for all methods, including classical knockoffs. Further, the values of the standard deviations for the TPR of the knockoff filter in the first two scenarios are higher than those of the other methods. This happens because, across the $100$ simulated data sets, the knockoff filter TPR is often zero and sometimes quite high. Finally, we note that classical knockoff has a slightly higher FDR than the threshold $0.1$ when the covariates are not independent and the covariance matrix is estimated using graphical lasso. As shown in Barber et al. (2020), the classical knockoff filter does not control FDR exactly below $q$ when the covariance matrix used to build the knockoffs is not the true one but an estimate. In our case, the graphical lasso tends to produce a sparser than needed graph. This behavior also supports our choice of setting the parameter of the priors in the proposed method to obtain a generally dense graph to be used to generate the knockoffs. Results for the third scenario confirm that the proposed Bayesian Knockoff filter is also able to outperform the state-of-art methods in settings with larger dimensionality. As a matter of fact, it successfully controls FDR under $0.1$, unlike the lasso regression. Moreover, the TPR is on average comparable to the one of the classical knockoff filter, but it is more stable as it has a much lower standard deviation.
Table 1: Simulation study: FDR, TPR, MCC and F1-score, for the scenarios described in Section 3.1. Compared methods are the proposed model (BayesKnock), classical Knockoff filter with exact (Knockoff (Exact)) and estimated (Knockoff (Glasso)) covariance matrix, Lasso, spike-and-slab regression (Spike-Slab), the joint model proposed by Peterson et al. (2016) (Joint) and the non local prior for variable selection proposed by Johnson and Rossell (2012). Values reported are averages over $100$ replicated data sets, with standard deviations in parentheses.
@@
Non-local prior 0.999 (0.01) 0.996 (0.02) 0.352 (0.07)
As for the comparison with the state-of-art methods for variable selection, we can see that Lasso has a high TPR in every scenario but fails to achieve any explicit control over the FDR, even in the simplest scenario. Compared with Bayesian variable selection methods via spike-and-slab priors, our model offers comparable results in the first two scenarios, where the covariates are independent, with almost perfect results. In the third scenario, with dependent covariates, classical spike-and-slab loses the control of FDR, unlike our method, even though it manages to find more relevant predictors than our method, while non-local spike-and-slab still controls the FDR but at the cost of a much lower TPR than our method. In Table 1 we do not show results for the BFDR approach proposed by Newton et al. (2004). Indeed, BFDR selected exactly the same variables as spike-and-slab in the first two scenarios. In the third scenario, however, the PPIs were hovering around $0.5$ and the BFDR did not select any variable. As pointed out in the introduction section, Bayes FDR selection procedures rely on the estimates of the marginal PPIs, as obtained from the MCMC output, while our procedure employs an upper bound on the probabilities of non-inclusion. Therefore, our model appears to achieve good operating characteristics and to control the FDR even when the estimates of the marginal posterior probability of inclusion fail to select the true relevant variables. This is evident when comparing our method with the model proposed by Peterson et al. (2016) (Joint). In scenarios 1 and 2 the two methods have the same performances, while in scenario 3, even though MCC and F1 have similar values, the addition of the knockoffs allows to keep the FDR under the desired level, while the joint method has higher FDR and therefore not an explicit control over it.
@@
our model achieved very solid performances for graph selection across the scenarios. For example, the average Frobenius norm was $1.404$, $1.415$ and $9.425$ and the F1-score was $0.869$, $0.865$ and $0.99$ in Scenarios $1$, 22 and $3$, respectively. Note that the Frobenius norm is much higher in Scenario $3$ due to the different magnitude in the correlation between variables. However, as we can see from the F1-score, the graph recovery still works.
We also show results on a simulation from an AFT model for survival outcome, with censored observations. We generate the covariates following the scheme employed in Scenario $3$ above, and then generate the log of the survival times from the linear AFT model. We allow for $25$ % of the subjects to be censored. Table 2 shows performances of our proposed method, compared with the spike-and-slab method for AFT models of Sha et al. (2004), the same AFT model without the latent knockoff layer , as also implemented by Peterson et al. (2016) and the non-local spike-and-slab prior as proposed by Johnson and Rossell (2012). These results are consistent with those for Scenario 3 in the previous simulation and also show that the proposed procedure suffers the least from the presence of censored data.
Table 2: Simulation study: FDR, TPR, MCC and F1-score, for the AFT models. Compared methods are the proposed model (BayesKnock), spike-and-slab regression (Spike-Slab), the joint model proposed by Peterson et al. (2016) (Joint) and the non local prior for variable selection proposed by Johnson and Rossell (2012). Values reported are averages over $100$ replicated data sets, with standard deviations in parentheses.